A ball has a momentum of 35 kg·m/s. If the mass of the ball is 5.0 kg, what is the speed of the ball?
1
Expert's answer
2020-02-24T10:20:27-0500

We can find the speed of the ball from the definition of the momentum:


"p = mv,""v = \dfrac{p}{m} = \dfrac{35 kg \cdot \dfrac{m}{s}}{5kg} = 7 \dfrac{m}{s}."

Answer:

"v = 7 \dfrac{m}{s}."